OpenStudy (praxer):

Find the Cubic in x which vanishes when x=1 and x=−2 and has values 4 and 8 when x=−1 and x=2 resprectively.

ganeshie8 (ganeshie8):

Let the required cubic be \(f(x) = k(x-1)(x+2)(x-a)\) plugin the given values and solve \(a,k\)

OpenStudy (praxer):

thank u : $$\frac{4x^3+2x^2-10x+4}{3}$$ and $$k=4/3\ and \ a=1/2$$